Experimental yield is 29 volume of hydrogen gas collected

Temperature of water/gas = 22 Celsius

Atmospheric pressure 765.8 mm Hg

1Mg+2HCI —> 1MgCl2+1H2

1. Using the balance chemical equation, solve for the volume of H2 that theoretically should be produced

2. Determine the percent yield by using the experimental yield/theoretical yield x 100
1
Expert's answer
2019-03-22T05:48:13-0400

m(Mg) = 36 g.

n(Mg) = m(Mg) / M(Mg) = 36 g / 24 g/mol = 1.5 mol.

1.5mol                           x mol

   Mg+2HCI —> MgCl2+H2

 1mol                             1mol

n(H2) = x = 1.5 mol.

p(H2)×V(H2)=n(H2)×R×T(H2)

R = 62.4 L×mm Hg / mol×K

TK = tC + 273.15 = (22+273.15) K = 295.15 K.

1)     V(H2)theor = n(H2)×R×T(H2)/ p(H2) = 1.5 mol × 62.4 L×mm Hg / mol×K × 295.15 K / 765.8 mm Hg = 36.07 L.

2)     η = V(H2)exper / V(H2)theor = 29 L / 36.07 L × 100% = 80.3992% ≈ 80.4%.